OpenStudy (anonymous):

If the fifth term of an arithmetic sequence is −5 and the ninth term is −17, the first term of this sequence must be? Would it be 7?

OpenStudy (kc_kennylau):

The \(n\)th term of an arithmetic sequence is \(a+d\times(n-1)\), where \(a\) is the first term and \(d\) is the common difference

OpenStudy (kc_kennylau):

Try to set up two equations from that :)

OpenStudy (anonymous):

how would I do that if I don't have the first term nor the second term to find the common difference?

OpenStudy (kc_kennylau):

You set up two equations to find that

OpenStudy (kc_kennylau):

The first equation is \(a+d\times(5-1)=-5\)

OpenStudy (kc_kennylau):

Can you figure out the second equation by yourself? :)

OpenStudy (nikato):

Sooty to intrude. But I got 7 too

OpenStudy (anonymous):

a+d X ( 9-1) = -17?

OpenStudy (anonymous):

oh really? I was just guessing on that haha

OpenStudy (kc_kennylau):

exactly

OpenStudy (kc_kennylau):

now you have a system of equations: \[\left\{\begin{array}{lr}a+4d=-5&---(1)\\a+8d=-17&---(2)\end{array}\right.\]

OpenStudy (anonymous):

ohhh not this thing haha ok

OpenStudy (anonymous):

I got 7?

OpenStudy (kc_kennylau):

yep :) \[2\times(1)-(2):a=7\]

OpenStudy (anonymous):

Thank you. You are the best

OpenStudy (kc_kennylau):

Thanks <3